OpenStudy (anonymous):

Please help! Solve the following applied problem using a system of equations. The sum of two numbers is 56. If the smaller number is 1/3 the larger number, what are the two numbers? If x is the smaller number, one equation is: a. x-y=56 b. x y=56 c. x/y=56 d. x+y=56

OpenStudy (zepp):

So sum of these two number in question is 56; x+y=56 this will be your first equation.

OpenStudy (zepp):

If the smaller number (x) is 1/3 the largest number, we have 1/3y = x

OpenStudy (anonymous):

Oh I see. So, c.? Can you please confirm?

OpenStudy (zepp):

Should be D)

OpenStudy (zepp):

\(x+y=56 \\x=\frac{1}{3}y\)

OpenStudy (anonymous):

Oh I get it now.

OpenStudy (anonymous):

Thank you zepp!

OpenStudy (zepp):

You are welcome :)